OpenStudy (anonymous):

A rectangular counter 3 feet wide and 5 feet long has a circle cut out of it in order to have a sink installed. The circle has a diameter of 18 inches. What is the approximate area of the remaining countertop surface? Use 3.14 for pi.

OpenStudy (anonymous):

Area of a rectangle is (length)(width)=3*5 square feet Area of a circle is (pi)(r)^2=(pi)(d/2)^2, but we should use same units. 18 inches = 1.5 feet Area of a circle is (3.14)(1.5/2)^2 square feet Subtract area of circle from area of rectangle.

OpenStudy (anonymous):

i got 4.71

OpenStudy (anonymous):

I did not get that.

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

Much bigger... this is a large counter top with a pretty small hole in it. What did you get for the rectangle? What did you get for the circle?

OpenStudy (anonymous):

2 pi r for the circle rectangle thats suppose to be 2l+2w

OpenStudy (anonymous):

Those are perimeter formulas... you want area. A=L*W for rectangle A=pi*r^2 for the circle

OpenStudy (anonymous):

half of 18 is 9 so that's the diameter

OpenStudy (anonymous):

That is another difficulty in this problem... the rectangle is described in feet. The circle is described in inches. You must choose which unit you need to work with. The radius of the circle is 9 inches. I chose to work in feet, but if you want to use inches, you must convert 3 feet and 5 feet to inches.

OpenStudy (anonymous):

3 times 5 which is 15

OpenStudy (anonymous):

Yes. The rectangle is 15 square feet.

OpenStudy (anonymous):

By saying that, you are choosing to work in feet. Convert the inches that you were given for the circle to feet by dividing by 12 inches. 18/12=1.5 feet for the diameter.

OpenStudy (anonymous):

To use the area equation for the circle, you will need radius. Radius is half of the diameter. Half of 1.5 feet is 0.75 feet. Now plug that number into A=pi*r^2 or pi*r*r

OpenStudy (anonymous):

i got 1.76625 and thats not one of my answers

OpenStudy (anonymous):

That is the area of the circle. Awesome! Now subtract that number from the size of the counter... 15-1.76625=?

OpenStudy (anonymous):

i got 13.23375

OpenStudy (anonymous):

Fantastic! That is the area of the remaining part of the counter in square feet.

OpenStudy (anonymous):

If your possible answers are much bigger, then they are probably in inches. Otherwise, we have the answer.

OpenStudy (anonymous):

yes thats one of the answers

OpenStudy (anonymous):

Yay! :)

OpenStudy (anonymous):

A medal for my efforts would be very nice. :)

OpenStudy (anonymous):

Thank you... now its time for bed. Good night.